The height of an inverted cone as it is filled with sand is always four times the radius. The radius is increasing at the rate of 3 cm/s. Calculate the rate at which the volume is increasing at the instant the volume is 800 cm^3
1
Expert's answer
2020-03-09T12:58:54-0400

Let us assume that:

The radius of inverted cone be "r" cm. and height be "h" cm.

Since:

The height of the cone is always four times the radius.

So:

We can write:

"h = 4r"

We are given:

The increasing rate of radius is:

"\frac{dr}{dt} = 3" cm/s.

And:

The volume of the cone is:

"V = 800 \,\, cm^3"

Now:

We know that, the volume of a cone is:

"V = \frac{1}{3} \pi r^2 h \\
\,\,\,\,\,\,= \frac{4}{3} \pi r^3 \hspace{1 cm} \left[ \because h = 4r \right]"

Again, we can also write:

"800 = \frac{4}{3} \pi r^3 \hspace{1 cm} \left[ \because V = 800 \right] \\
\Rightarrow r^3 = \frac{600}{\pi}"

"\Rightarrow r = \left( \frac{600}{\pi} \right)^{\frac{1}{3}}"

Now:

Derivative of "V" with respect to time "t" is:

"\frac{dV}{dt} = \frac{4}{3} \pi \frac{d}{dt} (r^3)"

"\,\,\,\,\,\,\,\,\,= \frac{4}{3} \pi \left( 3r^2 \frac{dr}{dt} \right)"

"\,\,\,\,\,\,\,\,\,= 4 \pi \left(\frac{600}{\pi} \right)^{\frac{2}{3}} (3)" "\hspace{1cm} \left[ \because \frac{dr}{dt} = 3 \, and \, r = \left( \frac{600}{\pi} \right)^{\frac{1}{3}} \right]"

"\,\,\,\,\,\,\,\,\,= 12 (\pi)^{\frac{1}{3}} (600)^{\frac{2}{3}}"

"\,\,\,\,\,\,\,\,\,\approx 1250"

Hence:

The volume of the cone is increasing approximately at the rate of "1250" cubic centimeters per second.